Considering the rotation of the Earth only, compare your apparent weight on the North Pole and on the Equator.
Answer #1

Weight on equator is less. Because of centrifugal force. Net accn on equator is

a = g- Rw^2

while on poles it is exact g

hence

Wpole / Weq = g / (g-Rw^2)

= 9.83 / (9.83 - 6400000 x (7.3x10^-5)^2)

= 1.0035